OpenStudy (anonymous):

4x+5y=2 -4x+y=34 again i only need the correct answer i do not understand what or how to do it and i can not afford to gamble with my college degree

OpenStudy (anonymous):

(-7,6)

OpenStudy (gw2011):

x=-7 and y=6

OpenStudy (anonymous):

If you don't want to gamble with your college degree, Why are you on here asking for answers. Aren't you going to have an exam that is going to make you display the techniques anyways?

OpenStudy (anonymous):

outcast not your problem obviously just leave me alone ok im pregnant and do not need stress of those just looking to be immature or aruge thanks for your input though

OpenStudy (anonymous):

from -4x+y = 34, y=34 + 4x substitute y = 34 + 4x into 4x+5y=2 4x+5(34+4x)=2 4x+170+20x = 2 24x = -168 x= -7 substitute x= -7 into y=34+4x y=34+4(-7) y=34-28 y=6 x=-7 y=6